A radio costs £21.60 in a sale. If it had a reduction of 10%, what was its original price?
Answer by LinnW(1048)

Set x = the original price.
(x - (10%)x) = 21.60
(x - 0.10x) = 21.60
(0.90x) = 21.60
divide each side by 0.90
x = 21.60/0.90
x = 24
The original price was £24 .